OpenStudy (anonymous):

There are 36 members of the track team. 1/3 of the members are girls. How many on the team are NOT girls?

OpenStudy (anonymous):

hi again :D

OpenStudy (anonymous):

if 1/3 was on the team were girls. that means 2/3 werenot. 2/3 * 36 = 24 people on the team were not girls

OpenStudy (callisto):

number of girls = 36 *(1/3) = 12 no of boys = total no - no of girls = 36 -12 = work it out yourself

OpenStudy (anonymous):

sweet i got it right!

OpenStudy (anonymous):

great! this sheet of paper i am doing is a grade and i am getting all of these right, and i am understanding them, but agin its a grade so i better check it and since my tutor isnt available to help you guys are all i got! Thank you so much!

OpenStudy (anonymous):

no problem :D we are glad to help!

OpenStudy (callisto):

very welcome :)